# Probing Heavily Obscured AGN in Major Galaxy Mergers Using the mm–X-ray Correlation

### v1

@@ I Introduction
Major galaxy mergers can simultaneously trigger intense accretion onto supermassive black holes (SMBHs; e.g., Hopkins et al. 2008; Treister et al. 2012) while obscuring this activity behind large amounts of gas and dust, both on nuclear (e.g., Ricci et al., 2017a; Koss et al., 2018; Ricci et al., 2021a) and galaxy-wide (Buchner and Bauer, 2017; Blecha et al., 2018) scales. These critical events have also been linked to a high incidence of dual or closely separated AGN (e.g., Imanishi, 2009; Pfeifle et al., 2019b, a; Farrah et al., 2022). As a result, Active Galactic Nuclei (AGNs, i.e., rapidly accreting SMBHs) in late-stage mergers are often enshrouded in material that can reach Compton-thick ( cm-2; e.g., Koss et al. 2016; Yamada et al. 2021; Ricci et al. 2021b) levels, which hampers their detection and characterization. This obscuration creates a significant gap in the census of SMBH growth. A complete accounting of the dual AGN population is crucial for predicting gravitational wave event rates, enabling precise constraints on their number density and physical properties—measurements that will become feasible with next-generation space-based gravitational wave observatories (e.g., Nandra et al., 2013; Amaro-Seoane et al., 2017; Piro et al., 2023). Probing heavily obscured SMBH growth generally requires multi-wavelength observations (e.g., Koss et al., 2012). In major galaxy mergers, however, intense star formation further complicates the AGN identification at optical, infrared, and radio wavelengths. For example, Koss et al. (2011) reported a dual AGN detected in hard X-rays but missed in optical and radio observations, highlighting these difficulties. Optical emission-line diagnostics are also known to be unreliable in mergers due to extinction and dilution by star formation (Koss et al., 2010). Millimeter (mm) wavelengths offer a promising avenue to study AGN in such environments, as they are largely unaffected by dust, even at very high column densities ( cm-2; Hildebrand 1983). Detailed SED decomposition analyses (e.g., Efstathiou et al., 2022; Varnava et al., 2025) have shown that many (U)LIRGs harbor deeply buried AGN, further motivating the use of dust-insensitive tracers such as mm emission.
@@
(Ultra-)Luminous Infrared Galaxies or (U)LIRGs (Sanders et al., 1988), provide an ideal laboratory to test this correlation. Defined by their infrared luminosities ( for LIRGs and for ULIRGs), these sources often host galaxy mergers with both intense star formation and heavily-obscured AGN activity. Of particular interest is the dual AGN phase, in which late-stage mergers host two active nuclei separated by 10 kpc, frequently accompanied by extreme obscuration (Ricci et al., 2021a) that can almost exclusively be probed at mm wavelengths.
@@ II Sample and Observational Data
The ultimate goal of this study is to evaluate the feasibility of using high angular resolution observations from ALMA, in combination with the millimeter–X-ray relation, to identify potential hidden AGN pairs in close proximity. This search for nearby obscured AGN using high angular resolution ALMA data will be presented in an upcoming publication (Droguett-Callejas et al., in prep.). As an initial step, and taking advantage of the currently available data, we focus on a sample of (U)LIRGs. These sources are uniquely suited to probe the Compton-thick SMBH growth regimes and are prime environments for hosting heavily obscured (dual) AGN. We therefore searched the ALMA archive111https://almascience.nrao.edu/aq/ for observations of sources in the complete, volume-limited GOALS sample (Armus et al., 2009), which contains more than 200 (U)LIRGs drawn from the IRAS Revised Bright Galaxy Sample (RBGS; Sanders et al., 2003). Specifically, we focused on observations from ALMA programs 2016.2.00055.S and 2017.1.00767.S (PI: E. Treister), designed to detect 12CO(2-1) emission and signatures of potential outflows in the line wings. These programs provide deep continuum data and a sample with well-defined statistical properties, selected based on the following criteria: (1) , to ensure good visibility with both ALMA and the VLA for potential follow-up; and (2) CO(1–0) peak flux mK, to guarantee detection of the 12CO(2–1) line . Note that these requirements imply that there will be (U)LIRGs included in the Kawamuro et al. (2022) sample that will be automatically incorporated in this study, such as IRASF 05189-2524. The ALMA Band 6 (211–275 GHz) observations used in these projects provide an optimal balance of low optical depth, suitable spatial resolution, and data availability. To build our working sample, we further required sources with available X-ray detections in either the 2–10 keV or 14–150 keV bands. These criteria yield 11 systems observed with the Atacama Compact Array (ACA), with 5.4” angular resolution, which correspond to a representative spatial resolution of 3.4 kpc for the sources in the sample. For four of them, CGCG 436030, IRASF 143481447, IRASF 171381017, and IRASF 172070014, we also analyzed archival higher-resolution 12-m array data with a median angular resolution of 0.41”, corresponding to a typical spatial resolution of 0.33 kpc at the distance of these sources. While ACA observations have lower spatial resolution than the 1′′ data used by Kawamuro et al. (2022) for isolated AGN, they provide a valuable baseline for assessing how spatial resolution affects the detectability of obscured AGN.
@@ III.3.1 Star Formation Contribution to the X-ray emission
Ricci et al. (2021a) examined this same connection in their hard X-ray study of late-stage AGN mergers, using NuSTAR observations of (U)LIRGs from the GOALS sample. They compared the observed X-ray luminosities with those predicted from star formation alone, based on the SFR–X-ray relations of Ranalli et al. (2003) and Lehmer et al. (2010). Their results show that systems without clear AGN signatures have X-ray luminosities that are consistent with, or even below, the values expected from star formation. This agrees with the findings of Torres-Albà et al. (2018), who reported a flattening of the SFR–X-ray relation in (U)LIRGs, likely due to increased obscuration within their star-forming regions. In contrast, sources that host an AGN systematically display X-ray luminosities in excess of those predicted by star formation alone.
While this relation applies to the 2–10 keV band, we would also like to assess the expected star formation contribution at higher energies. Persic and Rephaeli (2007) showed that in powerful (U)LIRGs, the 2–10 keV emission is dominated by young HMXBs. Assuming that HMXBs dominate the SFR-driven X-ray output, we adopt the same photon index of (See Section III.2). Figure 1 presents the observed mm and X-ray values of our (U)LIRGs sample against their derived SFR, compared to different SFR-mm/X-ray correlations. We derived the SFR values for our sample of (U)LIRGs by following the expression given by Kennicutt (1998), corrected for a Kroupa (2001) Initial Mass Function (IMF) where . We derived the integrated IR luminosities [8-1000 µm] using the observed IR flux densities reported by Sanders et al. (2003) and the equations reported in Table 1 from Sanders and Mirabel (1996). Panels a) and b) show the observed X-ray values for our sample of (U)LIRGs in the 2-10 and 14-150 keV bands, respectively, against their derived SFR. From both of these panels, we can see that the Lehmer et al. (2010) relation predicts a higher contribution to the X-ray emission from star formation. In addition, our results are consistent with the study by Ricci et al. (2021a), which found that sources without a confirmed AGN also exhibit consistent or lower luminosities compared to the relation by Lehmer et al. (2010). Consistently, one of the sources confirmed to host an AGN by other AGN tracers, NGC 835, is found significantly above the relation.
@@ III.3.2 Star Formation Contribution to the mm emission
At millimeter wavelengths, Murphy et al. (2012) found a correlation between SFR and millimeter emission at 33 GHz using a sample of extranuclear and nuclear star formation regions in nearby galaxies. This relation uses SFRs computed using a Kroupa (2001) IMF and considers both a free-free emission and a synchrotron emission component (see equation 10 in Murphy et al. (2012)). We use the explicit frequency dependence of this correlation to extrapolate from 33 GHz, the frequency at which it was derived, to 230 GHz, the representative frequency of the ALMA band 6 observations. This is done considering a non-thermal radio index of = 0.85, based on the average spectral index found for star-forming regions in NGC 6946. Yun and Carilli (2002) also described an SFR-mm correlation based on a sample of IR-selected dusty starburst galaxies. This correlation used a combination of free-free, synchrotron, and dust emission to model the mm emission, and the LIR to derive SFR using Kennicutt (1998) corrected by a Kroupa (2001) IMF. The entire radio-to-FIR continuum flux spectrum is given by:
@@ IV Results and Discussion
An alternative explanation for the low X-ray luminosities of these systems is that they may be intrinsically X-ray faint, a phenomenon previously reported in (U)LIRGs (Franceschini et al., 2003; Lonsdale et al., 2006; Ricci et al., 2016). Interestingly, a similar X-ray weakness is commonly observed in LoBAL quasars (Green and Mathur, 1996), which have been linked to (U)LIRGs, suggesting a possible evolutionary or physical connection between these populations. Still, the lack of additional AGN diagnostics argues against this scenario. For IRASF 17207-0014, the line-of-sight column densities from Ricci et al. (2021a) allow us to estimate possible intrinsic X-ray luminosities corrected for obscuration at cm-2 and cm-2, as presented in Table 4. In this case, the source could fall into the regime where the source has either a combination of star formation and AGN contribution, or it is mostly dominated by an obscured AGN (column densities of cm-2). However, no additional AGN evidence has been reported in the literature, and all available diagnostics indicate instead a starburst classification (Iwasawa et al., 2011). A few (U)LIRGs lie in the intermediate region, with X-ray luminosities exceeding those expected from star formation alone. We interpret this excess as likely due to AGN activity. Indeed, NGC 835 and IRASF 14348-1447 are confirmed AGN hosts based on multiple diagnostics (see Vardoulaki et al. (2015), Song et al. (2022) and references therein), while IRASF 17138-1017 shows evidence of an AGN only from its Chandra hardness ratio (Torres-Albà et al., 2018).
@@ IV.3 Obscuration effects
As discussed above, the mm continuum emission is expected to remain optically thin even in heavily obscured systems, whereas obscuration effects are more prominent in the X-ray regime. When column density measurements are available, we can correct the X-ray luminosities for obscuration. However, when is unavailable, we cannot estimate the true X-ray luminosity. Eight of the eleven GOALS sources lack column density estimates, which is expected given that in the low-counts regime, it is not feasible to carry out complex X-ray fitting to determine the presence of an AGN component. Their X-ray luminosities could therefore be underestimated; however, in most cases, there is no evidence for an AGN, and the emission is consistent with an origin due to star formation, for which extinction corrections are expected to be modest. These sources generally lie in the location expected for pure star-formation emission. Concealing an AGN in these systems would require extreme obscurations, with cm-2; values that are rare in the general AGN population (e.g., Peca et al., 2024), and/or intrinsically under-luminous X-ray emission. Nevertheless, X-ray studies of nearby (U)LIRGs in GOALS (Ricci et al., 2021a) suggest that a significant fraction of late-stage mergers could harbor heavily obscured, as-yet undetected AGN—underscoring the importance of calibrating an alternative mm AGN tracer. For the (U)LIRGs within the cyan triangle, showing a significant mm excess above the star-formation prediction, one could attempt to estimate the AGN luminosity required to account for this excess, if a reliable determination of the dominant star-formation component is feasible.
@@ V Conclusions
ALMA is a partnership of ESO (representing its member states), NSF (USA) and NINS (Japan), together with NRC (Canada), NSTC and ASIAA (Taiwan), and KASI (Republic of Korea), in cooperation with the Republic of Chile. The Joint ALMA Observatory is operated by ESO, AUI/NRAO and NAOJ. The National Radio Astronomy Observatory is a facility of the National Science Foundation operated under cooperative agreement by Associated Universities, Inc. This research has made use of the NASA/IPAC Extragalactic Database (NED), which is operated by the Jet Propulsion Laboratory, California Institute of Technology, under contract with the National Aeronautics and Space Administration.

### v3

@@ I Introduction
Major galaxy mergers can simultaneously trigger intense accretion onto supermassive black holes (SMBHs; e.g., Hopkins et al. 2008; Treister et al. 2012) while obscuring this activity behind large amounts of gas and dust, both on nuclear (e.g., Ricci et al., 2017a; Koss et al., 2018; Ricci et al., 2021) and galaxy-wide (Buchner and Bauer, 2017; Blecha et al., 2018) scales. These critical events have also been linked to a high incidence of dual or closely separated AGN (e.g., Imanishi, 2009; Pfeifle et al., 2019b, a; Farrah et al., 2022). As a result, Active Galactic Nuclei (AGNs, i.e., rapidly accreting SMBHs) in late-stage mergers are often enshrouded in material that can reach Compton-thick ( cm-2; e.g., Koss et al. 2016; Yamada et al. 2021; Ricci et al. 2021) levels, which hampers their detection and characterization. This obscuration creates a significant gap in the census of SMBH growth. A complete accounting of the dual AGN population is crucial for predicting gravitational wave event rates, enabling precise constraints on their number density and physical properties—measurements that will become feasible with next-generation space-based gravitational wave observatories (e.g., Nandra et al., 2013; Amaro-Seoane et al., 2017; Piro et al., 2023). Probing heavily obscured SMBH growth generally requires multi-wavelength observations (e.g., Koss et al., 2012). In major galaxy mergers, however, intense star formation further complicates the AGN identification at optical, infrared, and radio wavelengths. For example, Koss et al. (2011) reported a dual AGN detected in hard X-rays but missed in optical and radio observations, highlighting these difficulties. Optical emission-line diagnostics are also known to be unreliable in mergers due to extinction and dilution by star formation (Koss et al., 2010). Millimeter (mm) wavelengths offer a promising avenue to study AGN in such environments, as they are largely unaffected by dust, even at very high column densities ( cm-2; Hildebrand 1983). Detailed SED decomposition analyses (e.g., Efstathiou et al., 2022; Varnava et al., 2025) have shown that many (U)LIRGs harbor deeply buried AGN, further motivating the use of dust-insensitive tracers such as mm emission.
@@
(Ultra-)Luminous Infrared Galaxies or (U)LIRGs (Sanders et al., 1988), provide an ideal laboratory to test this correlation. Defined by their infrared luminosities ( for LIRGs and for ULIRGs), these sources often host galaxy mergers with both intense star formation and heavily-obscured AGN activity. Of particular interest is the dual AGN phase, in which late-stage mergers host two active nuclei separated by 10 kpc, frequently accompanied by extreme obscuration (Ricci et al., 2021) that can almost exclusively be probed at mm wavelengths.
@@ II Sample and Observational Data
The ultimate goal of this study is to evaluate the feasibility of using high angular resolution observations from ALMA, in combination with the millimeter–X-ray relation, to identify potential hidden AGN pairs in close proximity. This search for nearby obscured AGN using high angular resolution ALMA data will be presented in an upcoming publication (M. Droguett-Callejas et al., in prep.). As an initial step, and taking advantage of the currently available data, we focus on a sample of (U)LIRGs. These sources are uniquely suited to probe the Compton-thick SMBH growth regimes and are prime environments for hosting heavily obscured (dual) AGN. We therefore searched the ALMA archive111https://almascience.nrao.edu/aq/ for observations of sources in the complete, volume-limited GOALS sample (Armus et al., 2009), which contains more than 200 (U)LIRGs drawn from the IRAS Revised Bright Galaxy Sample (RBGS; Sanders et al., 2003). Specifically, we focused on observations from ALMA programs 2016.2.00055.S and 2017.1.00767.S (PI: E. Treister), designed to detect 12CO(2-1) emission and signatures of potential outflows in the line wings. These programs provide deep continuum data and a sample with well-defined statistical properties, selected based on the following criteria: (1) , to ensure good visibility with both ALMA and the VLA for potential follow-up; and (2) CO(1–0) peak flux mK, to guarantee detection of the 12CO(2–1) line . Note that these requirements imply that there will be (U)LIRGs included in the Kawamuro et al. (2022) sample that will be automatically incorporated in this study, such as IRASF 05189-2524. The ALMA Band 6 (211–275 GHz) observations used in these projects provide an optimal balance of low optical depth, suitable spatial resolution, and data availability. To build our working sample, we further required sources with available X-ray detections in either the 2–10 keV or 14–150 keV bands. These criteria yield 11 systems observed with the Atacama Compact Array (ACA), with 5.4” angular resolution, which correspond to a representative spatial resolution of 3.4 kpc for the sources in the sample. For four of them, CGCG 436030, IRASF 143481447, IRASF 171381017, and IRASF 172070014, we also analyzed archival higher-resolution 12-m array data with a median angular resolution of 0.41”, corresponding to a typical spatial resolution of 0.33 kpc at the distance of these sources. While ACA observations have lower spatial resolution than the 1′′ data used by Kawamuro et al. (2022) for isolated AGN, they provide a valuable baseline for assessing how spatial resolution affects the detectability of obscured AGN.
@@ III.3.1 Star Formation Contribution to the X-ray emission
Ricci et al. (2021) examined this same connection in their hard X-ray study of late-stage AGN mergers, using NuSTAR observations of (U)LIRGs from the GOALS sample. They compared the observed X-ray luminosities with those predicted from star formation alone, based on the SFR–X-ray relations of Ranalli et al. (2003) and Lehmer et al. (2010). Their results show that systems without clear AGN signatures have X-ray luminosities that are consistent with, or even below, the values expected from star formation. This agrees with the findings of Torres-Albà et al. (2018), who reported a flattening of the SFR–X-ray relation in (U)LIRGs, likely due to increased obscuration within their star-forming regions. In contrast, sources that host an AGN systematically display X-ray luminosities in excess of those predicted by star formation alone.
While this relation applies to the 2–10 keV band, we would also like to assess the expected star formation contribution at higher energies. Persic and Rephaeli (2007) showed that in powerful (U)LIRGs, the 2–10 keV emission is dominated by young HMXBs. Assuming that HMXBs dominate the SFR-driven X-ray output, we adopt the same photon index of (See Section III.2). Figure 1 presents the observed mm and X-ray values of our (U)LIRGs sample against their derived SFR, compared to different SFR-mm/X-ray correlations. We derived the SFR values for our sample of (U)LIRGs by following the expression given by Kennicutt (1998), corrected for a Kroupa (2001) Initial Mass Function (IMF) where . We derived the integrated IR luminosities [8-1000 µm] using the observed IR flux densities reported by Sanders et al. (2003) and the equations reported in Table 1 from Sanders and Mirabel (1996). Panels a) and b) show the observed X-ray values for our sample of (U)LIRGs in the 2-10 and 14-150 keV bands, respectively, against their derived SFR. From both of these panels, we can see that the Lehmer et al. (2010) relation predicts a higher contribution to the X-ray emission from star formation. In addition, our results are consistent with the study by Ricci et al. (2021), which found that sources without a confirmed AGN also exhibit consistent or lower luminosities compared to the relation by Lehmer et al. (2010). Consistently, one of the sources confirmed to host an AGN by other AGN tracers, NGC 835, is found significantly above the relation.
@@ III.3.2 Star Formation Contribution to the mm emission
At millimeter wavelengths, Murphy et al. (2012) found a correlation between SFR and millimeter emission at 33 GHz using a sample of extranuclear and nuclear star formation regions in nearby galaxies. This relation uses SFRs computed using a Kroupa (2001) IMF and considers both a free-free emission and a synchrotron emission component (see equation 10 in Murphy et al. (2012)). We use the explicit frequency dependence of this correlation to extrapolate from 33 GHz, the frequency at which it was derived, to 230 GHz, the representative frequency of the ALMA band 6 observations. This is done considering a non-thermal radio index of = 0.85, based on the average spectral index found for star-forming regions in NGC 6946. Yun and Carilli (2002) also described an SFR-mm correlation based on a sample of IR-selected dusty starburst galaxies. This correlation used a combination of free-free, synchrotron, and dust emission to model the mm emission, and the to derive SFR using Kennicutt (1998) corrected by a Kroupa (2001) IMF. The entire radio-to-FIR continuum flux spectrum is given by:
@@ IV Results and Discussion
An alternative explanation for the low X-ray luminosities of these systems is that they may be intrinsically X-ray faint, a phenomenon previously reported in (U)LIRGs (Franceschini et al., 2003; Lonsdale et al., 2006; Ricci et al., 2016). Interestingly, a similar X-ray weakness is commonly observed in LoBAL quasars (Green and Mathur, 1996), which have been linked to (U)LIRGs, suggesting a possible evolutionary or physical connection between these populations. Still, the lack of additional AGN diagnostics argues against this scenario. For IRASF 17207-0014, the line-of-sight column densities from Ricci et al. (2021) allow us to estimate possible intrinsic X-ray luminosities corrected for obscuration at cm-2 and cm-2, as presented in Table 4. In this case, the source could fall into the regime where the source has either a combination of star formation and AGN contribution, or it is mostly dominated by an obscured AGN (column densities of cm-2). However, no additional AGN evidence has been reported in the literature, and all available diagnostics indicate instead a starburst classification (Iwasawa et al., 2011). A few (U)LIRGs lie in the intermediate region, with X-ray luminosities exceeding those expected from star formation alone. We interpret this excess as likely due to AGN activity. Indeed, NGC 835 and IRASF 14348-1447 are confirmed AGN hosts based on multiple diagnostics (see Vardoulaki et al. (2015), Song et al. (2022) and references therein), while IRASF 17138-1017 shows evidence of an AGN only from its Chandra hardness ratio (Torres-Albà et al., 2018).
@@ IV.3 Obscuration effects
As discussed above, the mm continuum emission is expected to remain optically thin even in heavily obscured systems, whereas obscuration effects are more prominent in the X-ray regime. When column density measurements are available, we can correct the X-ray luminosities for obscuration. However, when is unavailable, we cannot estimate the true X-ray luminosity. Eight of the eleven GOALS sources lack column density estimates, which is expected given that in the low-counts regime, it is not feasible to carry out complex X-ray fitting to determine the presence of an AGN component. Their X-ray luminosities could therefore be underestimated; however, in most cases, there is no evidence for an AGN, and the emission is consistent with an origin due to star formation, for which extinction corrections are expected to be modest. These sources generally lie in the location expected for pure star-formation emission. Concealing an AGN in these systems would require extreme obscurations, with cm-2; values that are rare in the general AGN population (e.g., Peca et al., 2024), and/or intrinsically under-luminous X-ray emission. Nevertheless, X-ray studies of nearby (U)LIRGs in GOALS (Ricci et al., 2021) suggest that a significant fraction of late-stage mergers could harbor heavily obscured, as-yet undetected AGN—underscoring the importance of calibrating an alternative mm AGN tracer. For the (U)LIRGs within the cyan triangle, showing a significant mm excess above the star-formation prediction, one could attempt to estimate the AGN luminosity required to account for this excess, if a reliable determination of the dominant star-formation component is feasible.
@@ V Conclusions
ALMA is a partnership of ESO (representing its member states), NSF (USA) and NINS (Japan), together with NRC (Canada), NSTC and ASIAA (Taiwan), and KASI (Republic of Korea), in cooperation with the Republic of Chile. The Joint ALMA Observatory is operated by ESO, AUI/NRAO and NAOJ. The National Radio Astronomy Observatory and Green Bank Observatory are facilities of the U.S. National Science Foundation operated under cooperative agreement by Associated Universities, Inc. This research has made use of the NASA/IPAC Extragalactic Database (NED), which is operated by the Jet Propulsion Laboratory, California Institute of Technology, under contract with the National Aeronautics and Space Administration.